​(Efficiency analysis)  The Brenmar Sales Company had a gross profit margin​ (gross profitsdivided by÷​sales) of 32 percent and sales of $9.8 million last year.  79 percent of the​ firm's sales are on​ credit, and the remainder are cash sales. ​ Brenmar's current assets equal $1.7 ​million, its current liabilities equal $297,000​, and it has $107,400 in cash plus marketable securities.

a. If​ Brenmar's accounts receivable equal $562,100​, what is its average collection​ period?

b. If Brenmar reduces its average collection period to 20 ​days, what will be its new level of accounts​ receivable?

c. ​Brenmar's inventory turnover ratio is 9.9 times. What is the level of​ Brenmar's inventories?
Answer #1

1.
=Accounts Receivable*365/(Sales*% Credit)
=562100*365/(9.8*10^6*79%)
=26.50045

2.
=Sales*%Credit*Average Collection Period/365
=9.8*10^6*79%*20/365
=424219.17808

3.
=Sales*(1-Gross Profit Margin)/Inventory Turnover
=9.8*10^6*(1-32%)/9.9
=673131.31313